OpenStudy (anonymous):

I need help with a problem. Calculate the number of possibilities described in the following scenario: A two-player Football game requires that one player select an offensive play and the other chooses a defensive play. The Offensive player first chooses from 29 different plays. (There are 15 Run plays, 12 Pass plays, and 2 Kick Plays.) The Defensive Player then chooses one of 18 different plays: (There are 10 cover plays and 8 Blitz plays.) What is the total number of match ups of Offensive and Defensive plays?

OpenStudy (loser66):

Is it not that \(\left(\begin{matrix}29\\1\end{matrix}\right)*\left(\begin{matrix}18\\1\end{matrix}\right)\)

OpenStudy (anonymous):

I think so, thank you for your help.